# ESPnet-SpeechLM: An Open Speech Language Model Toolkit

### v1

@@ Abstract
Jinchuan Tian1 Jiatong Shi1 William Chen1 Siddhant Arora1 Yoshiki Masuyama2Takashi Maekaku3Yihan Wu1,4Junyi Peng1,5 Shikhar Bharadwaj1Yiwen Zhao1Samuele Cornell1Yifan Peng1 Xiang Yue1Chao-Han Huck Yang6Graham Neubig1Shinji Watanabe1 1Carnegie Mellon University 2Mitsubishi Electric Research Laboratories 3LY Corporation 4Renmin University of China 5Brno University of Technology 6NVIDIA Correspondence: jinchuat@andrew.cmu.edu
@@ 4.1 Experimental Setups
Evaluation: Following Maiti et al. (2024); Tian et al. (2024b), we test word error rate (WER) for ASR; ASR WER, Speaker Similarty and Proxy MOS for TTS; sWUGGY and sBLIMP Nguyen et al. (2020) for AudioLM. We measure the TextLM ability using popular metrics like MMLU Hendrycks et al. (2020), ARC-Challenge (ARC-C) Clark et al. (2018), HellaSwag (HS)Zellers et al. (2019), and OpenBookQA (OBQA) Mihaylov et al. (2018).

### v2

@@ Abstract
Jinchuan Tian1 Jiatong Shi1 William Chen1 Siddhant Arora1 Yoshiki Masuyama2Takashi Maekaku3Yihan Wu1,4Junyi Peng1,5 Shikhar Bharadwaj1Yiwen Zhao1Samuele Cornell1Yifan Peng1 Xiang Yue1Chao-Han Huck Yang6Graham Neubig1Shinji Watanabe1 1Carnegie Mellon University 2Mitsubishi Electric Research Laboratories 3LY Corporation 4Renmin University of China 5Brno University of Technology 6NVIDIA Research Correspondence: jinchuat@andrew.cmu.edu
@@ 4.1 Experimental Setups
Evaluation: Following Maiti et al. (2024); Tian et al. (2024b), we test word error rate (WER) for ASR; ASR WER, Speaker Similarty and Proxy MOS for TTS; perplexity for AudioLM. We measure the TextLM ability using popular metrics like MMLU Hendrycks et al. (2020), ARC-Challenge (ARC-C) Clark et al. (2018), HellaSwag (HS)Zellers et al. (2019), and OpenBookQA (OBQA) Mihaylov et al. (2018).